OpenStudy (anonymous):

The second derivative of g(x)=x^2-27/x(x>0) changes sign at the point x= ?

OpenStudy (anonymous):

can u use the equation editor so that makes more sense?

myininaya (myininaya):

is it \[g(x)=\frac{x^2-27}{x} or g(x)=x^2-\frac{27}{x}?\]

OpenStudy (anonymous):

sorry.The second one is correct.

OpenStudy (anonymous):

\[g'(x)=2x + \frac{27}{x^2}\]

myininaya (myininaya):

g'(x)=2x+(27/x^2 ) is this what you got for the fist derivative? then we need to find g'' g''(x)=2-(2*27/x^3)

OpenStudy (anonymous):

\[g''(x)=x-\frac{27*2}{x^3}\]

myininaya (myininaya):

now set g''=0 \[\frac{2x^3-2*27}{x^3}=0\] \[\frac{x^3-27}{x^3}=0\] \[\frac{(x-3)(x^2+3x+9)}{x^3}=0\] x=3

myininaya (myininaya):

cjn, (2x)'=2 not x

OpenStudy (anonymous):

yeah i saw that, thanks. disconnect between brain and keyboard.

myininaya (myininaya):

any questions ry?

myininaya (myininaya):

you need to check if the sign changes at 3 by pluggin in number before and after 3 to see if the sign does change

myininaya (myininaya):

g''(1)=- g''(4)=+ so yes the sign does change at x=3

OpenStudy (anonymous):

I see.i put g'(x)=0 in mistake.Thanks anyway!!

myininaya (myininaya):

? i thought you were looking for x not g'?

OpenStudy (anonymous):

i put g'(x)=0 instead of g''(x)=0 yea is finding x.

myininaya (myininaya):

ok

OpenStudy (anonymous):

thanks anyway:)